# Arena-Lite: Efficient and Reliable Large Language Model Evaluation via Tournament-Based Direct Comparisons

### v4

@@ Abstract
Seonil Son1††thanks: Current institution: RLWRLD, Ju-Min Oh1††thanks: Current institution: Samsung AI Research, Heegon Jin1††thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1 1NCSOFT Correspondence: seonil.calendar@gmail.com
@@ 1 Introduction
LLMs excel in diverse tasks, from chatbots to code generation, due to their powerful generative capabilities Ouyang et al. (2022); Roziere et al. (2023). As their versatility grows, accurately evaluating their performance becomes critical. To address this, benchmarks like MMLU and BigBench have emerged to assess LLM capabilities across various domains Hendrycks et al. (2020); Srivastava et al. (2023). Many of these benchmarks, such as those for arithmetic or code execution (e.g., GSM-Hard, HumanEval Gao et al. (2022); Chen et al. (2021)), use automated scoring to evaluate problem-solving skills. However, their focus is not on quality of generated content or limited to the cases where the generated contents are automatically evaluated (e.g. programming), which are mostly not the case for variety of generation tasks. The Chatbot Arena, a leading platform for reliable human evaluation of LLMs, has set a standard by collecting extensive human annotations Chiang et al. (2024). Yet, its resource-intensive approach has prompted efforts to replicate its rankings using LLM judges as a cost-effective alternative Li et al. (2024, 2023). These methods, however, rely on baseline-mediated comparisons—comparing LLM outputs to a baseline encore model’s outputs—which sacrifice reliability.
Current benchmarks relying on encore models often rank LLMs by their win rate against baseline responses from an encore model. This approach has two advantages: it scales linearly with the number of LLMs and provides a consistent quality standard. However, we argue that comparing LLMs directly against each other is inherently more reliable than using baseline outputs, which can introduce noise coming from weak transitivity Xu et al. (2025) of human preferences on LLM responses. To address this, we propose Arena-Lite, a novel evaluation framework that uses direct, head-to-head comparisons organized in a tournament structure. By eliminating the need for baseline outputs, Arena-Lite reduces the number of comparisons required while achieving stronger alignment with human-established rankings, such as those from Chatbot Arena.
@@ 2.1 Measuring Win rate over baseline outputs
Benchmarks like AlpacaEval and Arena-Hard-Auto assess LLM response quality by comparing it to baseline responses from a encore model Li et al. (2023, 2024). An LLM judge evaluates whether the candidate LLM’s response outperforms the baseline for a given prompt. The win rate—the proportion of prompts where the LLM’s response is preferred—serves as a measure of its generative ability. While this approach is straightforward and scalable, it introduces noise coming from mediated comparisons.
@@ 7 Conclusion
We introduced Arena-Lite, an efficient and reliable framework for evaluating Large Language Models (LLMs) through tournament-based direct comparisons. By eliminating the need for baseline encore outputs and adopting head-to-head comparison, Arena-Lite achieves higher reliability in system rankings with reduced number of comparisons. Our experiments, encompassing controlled stochastic modeling and empirical validation with various LLM judges, confirm that Arena-Lite consistently outperforms standard baseline-mediated evaluation methods, even with smaller datasets or weaker judges. The release of an accessible web demonstration and code supports the adoption of Arena-Lite to help streamlining model development cycle across research and industry. Future work will extend Arena-Lite’s application to diverse domains, including multi-modal LLM evaluation involving visual or audio inputs and outputs.
@@ A.2 Full table for Experiment 2
Here is the extended results of Experiment 2 (Section 4.3) presented in Table 3. Aligned LLMs smaller than 7B parameters struggles to work as a proper Judge. Otherwise, Arena-Lite method excels over common practice of using encore outputs as baselines.
| Dataset size | method | claude 3.5 sonnet | llama3.1-8b-it | qwen2.5-7b-it | qwen2.5-0.5b-it | gemma2-2b-it |

### v5

@@ Abstract
Seonil Son1††thanks: Current institution: RLWRLD Inc., Ju-Min Oh1††thanks: Current institution: Samsung AI Research, Heegon Jin1††thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1††thanks: Current institution: Samsung Life Insurance 1NC AI Correspondence: seonil.calendar@gmail.com
@@ 1 Introduction
LLMs excel in diverse tasks, from chatbots to code generation, due to their powerful generative capabilities Ouyang et al. (2022); Roziere et al. (2023). As their versatility grows, accurately evaluating their performance becomes critical. To address this, benchmarks like MMLU and BigBench have emerged to assess LLM capabilities across various domains Hendrycks et al. (2020); Srivastava et al. (2023). Many of these benchmarks, such as those for arithmetic or code execution (e.g., GSM-Hard, HumanEval Gao et al. (2022); Chen et al. (2021)), use automated scoring to evaluate problem-solving skills. However, their focus is not on quality of generated content or limited to the cases where the generated contents are automatically evaluated (e.g. programming), which are mostly not the case for variety of generation tasks. The Chatbot Arena, a leading platform for reliable human evaluation of LLMs, has set a standard by collecting extensive human annotations Chiang et al. (2024). Yet, its resource-intensive approach has prompted efforts to replicate its rankings using LLM judges as a cost-effective alternative Li et al. (2024, 2023). These methods, however, rely on baseline-mediated comparisons—comparing LLM outputs to a leading proprietary model’s outputs—which sacrifice reliability.
Current benchmarks relying on baseline often rank LLMs by their win rate against baseline responses from an leading proprietary models. This approach has two advantages: it scales linearly with the number of LLMs and provides a consistent quality standard. However, we argue that comparing LLMs directly against each other is inherently more reliable than using baseline outputs, which can introduce noise coming from weak transitivity Xu et al. (2025) of human preferences on LLM responses. To address this, we propose Arena-Lite, a novel evaluation framework that uses direct, head-to-head comparisons organized in a tournament structure. By eliminating the need for baseline outputs, Arena-Lite reduces the number of comparisons required while achieving stronger alignment with human-established rankings, such as those from Chatbot Arena.
@@ 2.1 Measuring Win rate over baseline outputs
Benchmarks like AlpacaEval and Arena-Hard-Auto assess LLM response quality by comparing it to baseline responses from proprietary model Li et al. (2023, 2024). An LLM judge evaluates whether the candidate LLM’s response outperforms the baseline for a given prompt. The win rate—the proportion of prompts where the LLM’s response is preferred—serves as a measure of its generative ability. While this approach is straightforward and scalable, it introduces noise coming from mediated comparisons.
@@ 7 Conclusion
We introduced Arena-Lite, an efficient and reliable framework for evaluating Large Language Models (LLMs) through tournament-based direct comparisons. By eliminating the need for baseline outputs and adopting head-to-head comparison, Arena-Lite achieves higher reliability in system rankings with reduced number of comparisons. Our experiments, encompassing controlled stochastic modeling and empirical validation with various LLM judges, confirm that Arena-Lite consistently outperforms standard baseline-mediated evaluation methods, even with smaller datasets or weaker judges. The release of an accessible web demonstration and code supports the adoption of Arena-Lite to help streamlining model development cycle across research and industry. Future work will extend Arena-Lite’s application to diverse domains, including multi-modal LLM evaluation involving visual or audio inputs and outputs.
@@ A.2 Full table for Experiment 2
Here is the extended results of Experiment 2 (Section 4.3) presented in Table 3. Aligned LLMs smaller than 7B parameters struggles to work as a proper Judge. Otherwise, Arena-Lite method excels over common practice of using outputs from proprietary as baselines.
| Dataset size | method | claude 3.5 | llama3.1 | qwen2.5 | qwen2.5 | gemma2 |
| sonnet | 8b-it | 7b-it | 0.5b-it | 2b-it | ||
@@ A.8.2 Decoding Parameters
] # prompt ends here
LLMBar prompt of our use. We used metric variant suggested in original LLMBar paper. More preset prompts are in our Arena-Lite Demo and source (https://huggingface.co/spaces/NCSOFT/ArenaLite)